There is a claim that on daily average students spend 5.5 hours on social media with standard deviation of 1.25 .A resercher wants to test the claim at 5% level significance.



1.formulate the null and alternatuve hyphothesis


2.Determine the test statistic to be used


3.Find the corresponding z value


4. Identify the rejection region


1
Expert's answer
2022-06-08T16:18:05-0400

1. The following null and alternative hypotheses need to be tested:

"H_0:\mu=5.5"

"H_1:\mu\not=5.5"

2. This corresponds to a two-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

3. Based on the information provided, the significance level is "\alpha = 0.05," and the critical value for a two-tailed test is "z_c = 1.96."

4. The rejection region for this two-tailed test is "R = \{z:|z|>1.96\}."


The z-statistic is computed as follows:



"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}"